You have two flashlights that have 1.00-A currents flowing through them. One flashlight has a single 1.50-V battery in its circuit, and the second flashlight has three 1.50-V batteries connected in a chain that provides 4.50 V. How much power is the battery in the first flashlight providing? How much power is each battery in the second flashlight providing?
Answer #1

Power provided by battery is given by power = voltage* current= V*I

For first flashlight, voltage is 1.5 V and current is 1A.

So power is P= 1.5*1= 1.5 W

For second flashlight, voltage across flashlight is 4.5V and current is 1A. So total power provided by 3 identical batteries is P= 4.5*1= 4.5 W. As the batteries are identical, power provided by each battery is 4.5/3= 1.5W